Describe how and why the multiplier and the interest sensitivity of aggregate demand affect the slope of the IS curve


1
Expert's answer
2021-09-24T11:29:29-0400

The slope of the IS curve is 1 / (kAb) or MLR / b, where MLR is the maximum seizure rate (recall that MLR = 1 - mpc (1 - t) + mpm = mps (1 - t) + t + mpm, i.e. That is, the marginal rate of withdrawals is the reciprocal of the multiplier of expenditures, MLR = 1 / kA). Thus, the slope of the IS curve is determined by: 1) the sensitivity of autonomous expenditures to the interest rate (b),


2) the value of the multiplier (kA), which depends on the marginal propensity to consume (mpc), the tax rate (t) and the marginal propensity to import (mpm).


The slope of the IS curve decreases (it turns clockwise and becomes flatter). The IS curve will be flatter:


The sensitivity of autonomous expenditures to the interest rate (b) is high, which


means that even a small change in the interest rate leads to a significant change in autonomous costs and, therefore, income;


The expense multiplier (kA) is large and the maximum withdrawal rate (MLR) is small,


what is possible if: a) the marginal propensity to consume is high; b) the marginal tax rate is low; c) the marginal propensity to import is small. If the multiplier is large, it means that even a small change in autonomous expenses will lead to a large multiplicative change in income. (Note that the magnitude of the multiplier determines both the slope and the magnitude of the shift of the IS curve).